# KineSoft: Learning Proprioceptive Manipulation Policies with Soft Robot Hands

### v1

@@ IV-B Shape Estimation Model
The objective of the shape estimation model is to learn a mapping between sensor readings from embedded strain sensors $\mathbf{S}$ and the vertex displacements of the MOE fingers. Each finger contains four strain sensors arranged between the tendons, giving us $n=12$ resistance measurements total for the three-fingered hand, denoted as $\mathbf{R}\in\mathbb{R}^{n}$.
The model learns a function $f$ that maps from sensor resistance readings and initial vertex positions to displacement vectors. Let $\mathbf{V}_{j,0}\in\mathbb{R}^{N\times 3}$ denote the initial vertex positions of finger $j$, where $N$ is the number of vertices in the mesh. The mapping can be expressed as:
$f:(\mathbf{R},\{\mathbf{V}_{j,0}\}_{j=1}^{3})\mapsto\{\Delta\mathbf{V}_{j}\}_{j=1}^{3}$
We implement this mapping using a FoldingNet-inspired architecture. The network first encodes each finger’s four strain measurements into a latent representation $\mathbf{z}_{j}\in\mathbb{R}^{128}$ through an encoding function:
$\mathbf{z}_{j}=h_{\text{enc}}(\mathbf{R}_{j})$
The decoder then processes each vertex independently by learning a mapping from the concatenated initial vertex position and latent displacement:
$\Delta\mathbf{v}_{j,i}=h_{\text{dec}}(\mathbf{v}_{j,0}^{i},\mathbf{z}_{j}),$
where $\mathbf{v}_{j,0}^{i}\in\mathbb{R}^{3}$ is the initial position of vertex $i$ in finger $j$, and [;][;][ ; ] denotes concatenation. This vertex-wise processing allows the network to learn local deformation patterns while maintaining spatial relationships defined by the initial mesh topology. The deformed vertex positions are then obtained by applying the predicted displacements:
$\mathbf{v}_{j,t}^{i}=\mathbf{v}_{j,0}^{i}+\Delta\mathbf{v}_{j,i}$
The FoldingNet decoder architecture enables learning vertex-level deformations while leveraging the spatial structure of the initial mesh configuration. By predicting displacements rather than absolute positions, the network learns naturally centered and scaled deformations, leading to more stable training and better generalization. To train the model, we generate a large dataset of deformed meshes using SOFA (Simulation Open Framework Architecture) [41]. Our simulation setup uses a tetrahedral finite element mesh of the MOE finger with a Neo-Hookean hyperelastic material model parameterized by elastic material properties that are randomized at runtime.
We model the tendon actuation with massless, inextensible cables running through a series of fixed points within the finger body. Each tendon path is discretized into segments defined by 3D attachment points embedded in the tetrahedral mesh. The cable constraint applies forces to these points to maintain constant length while allowing sliding, effectively simulating the mechanical behavior of Bowden cable transmission. The soft body scene is solved with an implicit Euler time integration scheme and uses a conjugate gradient solver for the system matrices. We generate training data by randomly sampling tendon actuation commands within the feasible range and recording the resulting deformed vertex positions and embedded sensor strains. To simulate rich deformation behaviors including contact-like effects, we apply random external forces to the finger surface. These forces are randomly applied over time with sufficiently large radii to ensure smooth deformations that mimic natural contact interactions, without requiring explicit and difficult-to-model contacts in the scene.
The model is trained using mean squared error (MSE) loss on vertex displacements:
where $\Delta\mathbf{v}_{j,i}^{*}$ represents the ground truth displacement for vertex $i$ of mesh finger $j$. This choice of loss function provides strong supervision by enforcing explicit vertex-wise correspondence between predicted and ground truth meshes. Because we leverage simulated data to train the model, we can exploit the vertex-level correspondences in the meshes unlike prior works that had to rely on chamfer distance loss over real-world partial observations [38], MSE loss ensures that each vertex learns to track its specific local deformation patterns, enabling precise reconstruction of the full finger shape.
IV-C Sim-to-Real Domain Alignment
@@
$\frac{L_{i,t}^{R}-L_{i,0}^{R}}{L_{i,0}^{R}}=\frac{1}{1+\frac{\epsilon_{i}}{L_{i,0}^{S}}}\cdot\frac{L_{i,t}^{S}-L_{i,0}^{S}}{L_{i,0}^{S}}$ (4)
@@
where the observed resistances $R_{i,t},R_{i,0}$ are measured with the DAQ setup. For the $n$ embedded sensors, aligning the simulated and observed distributions involves optimizing the constant correction parameters $\kappa_{0},\kappa_{1},\dots,\kappa_{n-1}$. The objective is defined as:
where $T$ is the number of time steps for which resistance data is observed.
Finally, because the corresponding simulated lengths for an observed resistance are not available (i.e., there is no direct sim-to-real correspondence between the observed resistance values and the simulated lengths), we instead optimize the correction factors $\kappa_{0},\kappa_{1},\dots,\kappa_{n-1}$ by minimizing the Chamfer distance between the observed points on the fingers and the model’s predicted surface points. This optimization assumes that there is a unique mapping between the internal strain of the sensors and the deformation of the soft structure, which is inferred through the observed resistance values.
The model defines a mapping from the sensor resistance observations to the surface mesh of the fingers based on the predicted displacements of the vertices. However, in the real-world deployment of the model as shown in the Figure 2, we do not have access to labeled pairs of real-world resistance values $\mathbf{R}$ and $\mathbf{L^{R}}$:
$\mathcal{L}_{\text{UCD}}=\sum_{j=0}^{m-1}\sum_{\mathbf{p}_{\text{obs}}\in\mathcal{P}_{\text{obs}}}\min_{\mathbf{v}\in\mathbf{V}_{j,t}}\|\mathbf{p}_{\text{obs}}-\mathbf{v}\|^{2},$ (7)
where $\mathcal{P}_{\text{obs}}$ is the set of observed points on the surface of the fingers from the RGB-D camera feed and $\mathbf{V}_{j,t}$ are vertices of the predicted surface mesh of finger $j$ defined by $\mathbf{M}_{j,t}=f(\mathbf{R},\mathbf{V}_{j,0})$. We compute $\mathcal{L}_{\text{UCD}}(\mathbf{\kappa})$ where the correction factors $\kappa_{i}$ are adjusted such that the predicted surface points align with the observed deformation of the fingers.
The function $f$ is differentiable and thus gradient descent-based optimization is possible here for locally optimal $\mathbf{\kappa}$. However, because of the noisy loss landscape and relatively cheap inference cost, we achieved better performance by optimizing $\mathbf{\kappa}$ with a sample-based evolutionary strategy as shown in Algorithm 1. By optimizing for the alignment factors $\mathbf{\kappa}$, the model effectively aligns the observed resistance-to-strain relationship with the real-world deformation states, ensuring robust sim-to-real transfer for the shape estimation model.
@@
IV-D Shape-conditioned Controller
The shape-conditioned controller leverages real-time proprioceptive mesh state estimation of the MOE fingers to track desired shape trajectories. For each finger $j$, the controller compares the current estimated vertex positions $\mathbf{V}_{t}$ with desired target positions $\mathbf{V}_{t}^{D}$ generated from the policy trajectory. The shape residual is calculated over the corresponding vertices of the meshes:
$\mathbf{e}_{j,t}=\mathbf{V}_{t,j}^{D}-\mathbf{V}_{t,j}$
Each finger is actuated by a pair of antagonistic tendons controlled by two servos. The actuation directions for each servo pair are represented by unit vectors $\mathbf{d}_{2j},\mathbf{d}_{2j+1}\in\mathbb{R}^{2}$ that capture the primary deformation modes. The servo adjustments $\delta u_{j,t}$ for each finger are computed by projecting the shape error onto these actuation directions:
$\delta u_{j,t}=k_{p}\sum_{n}\mathbf{e}_{j,t}^{n}\cdot[\mathbf{d}_{2j},\mathbf{d}_{2j+1}]^{T}$
where $k_{p}$ is a scalar gain and $\mathbf{e}_{j,t}^{n}$ is the error for vertex $n$. In deployment, the actions are clipped to prevent overloading the actuators. This controller runs at 100Hz with the shape estimation at each step, enabling responsive shape trajectory tracking. By projecting shape errors onto fitted actuation directions, the controller effectively translates desired deformations into appropriate servo commands despite the complex relationship between tendon actuation and finger deformation.
IV-E Imitation Policy
The shape estimation model from Section IV-B provides proprioceptive feedback through predicted vertex positions $\mathbf{V}_{t}$. The wrist-mounted RGB-D camera captures point cloud observations of the manipulation workspace. Using these complementary sources of state information, we train a diffusion policy to learn manipulation skills via imitation learning.
The policy predicts actions that combine shape deformations and end-effector pose changes:
$a_{t}=\{\Delta\mathbf{V}_{t},\Delta\mathbf{p}_{t}\}$
The state representation combines two different point-based encodings:
$s_{t}=\{h_{\text{shape}}(\mathbf{V}_{t}),h_{\text{pc}}(\mathcal{P}_{t}),\mathbf{p}_{t}\},$
where $h_{\text{shape}}$ processes the predicted vertex positions using MLPs that leverage vertex correspondence, $h_{\text{pc}}$ is a DP3 point cloud encoder that processes subsampled RGB-D observations $\mathcal{P}_{t}$, and $\mathbf{p}_{t}$ is the current end-effector pose.
The diffusion model iteratively denoises a Gaussian distribution into meaningful actions through a reverse process:
$a_{t-1}=\mu_{\theta}(a_{t},s_{t},t)+\sigma_{t}\mathbf{z},$
where $\mu_{\theta}$ is a learned denoising model and $\mathbf{z}\sim\mathcal{N}(0,\mathbf{I})$.
The trained policy works with the shape-conditioned controller from Section IV-D, which maps the policy’s deformation predictions to actuator commands. This hierarchical approach enables robust manipulation by combining learned high-level behaviors with precise low-level tracking.
Figure 5: Shape Estimation Fidelity Results. We compare the reconstructed shapes against baselines from literature and ablate with our model-based approach from Section IV-B.
V Experiments
V-A Baselines
Shape-estimation Baselines. For shape estimation, we compare with analytical and learning-based baselines:
  • Constant Curvature Model [9, 45]. Constant curvature model is a common representation for the continuum deformation behavior of soft robot that parametrizes the shape with a single curvature curve [51]. Typically, the independent parameters of the state of the robot are defined by $r_{curve}$ and $\theta_{curve}$. Assuming a constant length, $L_{curve}$ of the robot, we get the constraint:
    $L_{curve}=r_{curve}\theta_{curve}.$
    In typical applications, additional term $\phi_{curve}$ is introduced to represent the plane of bending [17]. We implemented this simplified representation for soft robot shape using the proposed strain model as outlined in Section IV-C and fitting $r_{curve}$ and $\theta_{curve}$ to the observed strains in each side of the curve. We transformed the cross-section boundary to the curve during the evaluation and measured the chamfer distance to the reference.

### v2

@@ 4.2 Shape Estimation Model
The shape estimation model maps strain sensor readings $\mathbf{R}\in\mathbb{R}^{n}$ to vertex displacements of the MOE fingers. Each of the three fingers contains four embedded strain sensors, giving $n=12$ total resistance values. The model learns a function $f$ that predicts per-vertex deformations conditioned on both the sensor input and the undeformed mesh:
$f(\mathbf{R},\{\mathbf{V}_{j,0}\}_{j=1}^{3})=\{\Delta\mathbf{V}_{j}\}_{j=1}^{3},$
where $\mathbf{V}_{j,0}\in\mathbb{R}^{N\times 3}$ denotes the initial vertex positions of finger $j$, and $\Delta\mathbf{V}_{j}$ are the predicted displacements. We implement $f$ using a FoldingNet-based architecture [48]. For each finger $j$, an encoder $h_{\text{enc}}$ transforms its four resistance values $\mathbf{R}_{j}$ into a latent code $\mathbf{z}_{j}\in\mathbb{R}^{128}$ as $\mathbf{z}_{j}=h_{\text{enc}}(\mathbf{R}_{j})$.
A decoder $h_{\text{dec}}$ then predicts the displacement of each vertex $\mathbf{v}_{j,0}^{i}$ by combining its initial position with $\mathbf{z}_{j}$:
$\Delta\mathbf{v}_{j,i}=h_{\text{dec}}(\mathbf{v}_{j,0}^{i},\mathbf{z}_{j}),\quad\mathbf{v}_{j,t}^{i}=\mathbf{v}_{j,0}^{i}+\Delta\mathbf{v}_{j,i}.$
This formulation allows the model to learn local deformation behavior conditioned on global strain input while preserving mesh topology. Predicting displacements rather than absolute positions improves stability and generalization as shown in comparison to baselines. The model is trained on a dataset of simulated deformations as described in the Appendix paired with corresponding simulated strain readings.
4.3 Sim-to-Real Domain Alignment
To bridge the gap between simulation and real-world deployment, we introduce a calibration procedure to align sensor readings from physical experiments with the simulated deformation model. Our goal is to determine a set of correction factors $\kappa_{0},\dots,\kappa_{n-1}$, one per sensor, that align the strain-based deformation estimates in simulation with real-world resistance measurements.
We derive the following objective, based on the sensor model described in the Appendix, to align simulated deformations with real-world sensor readings:
where $n$ is the number of sensors, $T$ is the number of sample frames, $R_{i,t}$ is the resistance measured from the $i$-th sensor at time step $t$, and $L_{i,t}^{S}$ is the corresponding simulated internal length of the $i$-th sensor at time $t$ The correction factors $\kappa_{i}$ are optimized to compensate for the domain mismatch between real and simulated sensor responses.
In practice, the true simulated lengths corresponding to each real-world resistance reading are unavailable, and the sensor-to-geometry mapping is not directly observable. Thus, we approximate this mapping by minimizing the unsupervised Chamfer distance between the observed 3D positions of points on the soft fingers and the predicted surface points from our shape estimation model:
$\mathcal{L}_{\text{UCD}}=\sum_{j=0}^{m-1}\sum_{\mathbf{p}_{\text{obs}}\in\mathcal{P}_{\text{obs}}^{(j)}}\min_{\mathbf{p}_{\text{pred}}\in\mathcal{P}_{\text{pred}}^{(j)}}\|\mathbf{p}_{\text{obs}}-\mathbf{p}_{\text{pred}}\|^{2},$ (2)
where $\mathcal{P}_{\text{obs}}^{(j)}$ and $\mathcal{P}_{\text{pred}}^{(j)}$ denote the observed and predicted point clouds for the $j$-th example.
This alignment procedure ensures that the internal deformation estimated from resistance measurements produces a surface mesh consistent with external geometry observations. It connects to the overall pipeline by enabling the trained model to generalize from simulation to physical deployment without requiring paired real-world deformation labels.
Figure 3: Demonstration and KineSoft Rollout for Bottle Unscrewing Task.
4.4 Shape-conditioned Controller
The shape-conditioned controller leverages real-time proprioceptive mesh state estimation of the MOE fingers to execute desired shape trajectories. For each finger $j$, the controller compares the current estimated vertex positions $\mathbf{V}_{t}$ with desired target positions $\mathbf{V}_{t}^{D}$ generated from the policy trajectory. Each finger is actuated by a pair of antagonistic tendons controlled by two servos. The actuation directions for each servo pair are represented by unit vectors $\mathbf{d}_{2j},\mathbf{d}_{2j+1}\in\mathbb{R}^{2}$ that capture the primary deformation modes. The servo adjustments $\delta u_{j,t}$ for each finger are computed by projecting the shape error onto these actuation directions:
$\delta u_{j,t}=k_{p}\sum_{n}\mathbf{e}_{j,t}^{n}\cdot[\mathbf{d}_{2j},\mathbf{d}_{2j+1}]^{T},$
where $k_{p}$ is a scalar gain and $\mathbf{e}_{j,t}^{n}$ is the shape error for vertex $n$ defined by $\mathbf{e}_{j,t}=\mathbf{V}_{t,j}^{D}-\mathbf{V}_{t,j}$. In deployment, the controller clips actions to prevent overloading the actuators. The controller runs at 100 Hz with the shape estimation at each step, enabling responsive shape trajectory tracking. By projecting shape errors onto fitted actuation directions, the controller effectively translates desired deformations into appropriate servo commands despite the complex relationship between tendon actuation and finger deformation.
4.5 Imitation Policy
The shape estimation model (Section 4.2) provides proprioceptive feedback via predicted surface vertex positions $\mathbf{V}_{t}$, while a wrist-mounted RGB-D camera captures external point cloud observations $\mathcal{P}_{t}$ of the workspace. We train a diffusion policy to imitate manipulation skills using these complementary inputs.
The policy predicts actions $a_{t}=\{\Delta\mathbf{V}_{t},\Delta\mathbf{p}_{t}\}$ that couple surface deformations and end-effector motion. The state $s_{t}=\{h_{\text{shape}}(\mathbf{V}_{t}),h_{\text{pc}}(\mathcal{P}_{t}),\mathbf{p}_{t}\}$ combines proprioceptive and exteroceptive information: $h_{\text{shape}}$ is an MLP encoder applied to mesh vertices with temporal correspondence, $h_{\text{pc}}$ is a DP3 encoder for the RGB-D point cloud, and $\mathbf{p}_{t}$ is the current end-effector pose.
The diffusion-based imitation policy learns to denoise a vertex-level action trajectory via the reverse process $a_{t-1}=\mu_{\theta}(a_{t},s_{t},t)+\sigma_{t}\mathbf{z}$, where $\mu_{\theta}$ is a learned denoising network and $\mathbf{z}\sim\mathcal{N}(0,\mathbf{I})$. Combined with the shape-conditioned controller (Section 4.4), the policy enables robust, contact-rich manipulation by fusing learned high-level intent with precise low-level tracking.
5 Experiments
We evaluate KineSoft through real-world experiments that assess each component of the framework and key design choices. We quantify the accuracy of the shape estimation model, followed by evaluating the shape-conditioned controller’s ability to track target deformations. Finally, we assess KineSoft policies across six in-hand manipulation tasks involving both rigid and deformable objects. These experiments demonstrate the effectiveness of shape-based representations for proprioceptive control and reliable manipulation, particularly in settings where soft robots offer inherent advantages.
5.1 Shape Estimation
Table 1: Shape Estimation Fidelity
Method Shape Error [mm]
PneuFlex Sensor $3.70\pm 0.01$.36)
DeepSoRo $3.27\pm 0.01$.05)
KineSoft (naive) $4.91\pm 0.02$.85)
KineSoft (unconstrained) $4.36\pm 0.03$.47)
KineSoft (model) $1.92\pm 0.00$.39)
We evaluated the proposed shape estimation model against baselines from the literature: the constant curvature model [49, 41], an analytical representation of soft robot deformation, and DeepSoRo [50], a learning-based point cloud reconstruction method. In addition, we compared against two ablations of our approach: a naively calibrated linear mapping and an unconstrained fully connected model, both trained using the domain alignment procedure described in Section 4.3. Evaluation was conducted using observed point clouds captured from the physical setup shown in Figure 7, and performance was measured using the unidirectional Chamfer distance defined in Equation 2. Quantitative results are summarized in Table 1 and visualized in Figure 6. Our model achieved a shape estimation error of 1.92,mm, representing a 41.3% improvement over the best baseline (DeepSoRo) and a 60.9% improvement over the linear variant of our method.
5.2 Shape-conditioned Controller Performance
Table 2: Shape Tracking Error Comparison
Method Representation Error [mm]
Strain-tracking Strain $6.20\pm 0.02$.39)
KineSoft Mesh $3.29\pm 0.00$.91)
We evaluated the proposed shape-conditioned controller against a strain-tracking baseline that directly uses sensor readings for control as implemented in prior soft robot manipulation works [3, 51]. For evaluation, we collected reference trajectories through kinesthetic teaching, where a demonstrator physically deformed the fingers into desired configurations. During execution, the controllers had to track these trajectories using tendon actuation. This evaluation highlights a fundamental challenge in soft robot imitation: the sensor signals generated during kinesthetic demonstration (when fingers are manually deformed) differ significantly from those produced during autonomous execution (when tendons are actuated). The strain-tracking baseline, which attempts to directly match these sensor readings, struggles with this demonstration-execution gap, achieving 6.20 mm tracking error. In contrast, the proposed shape-conditioned controller bridges this gap by tracking the geometric shape itself rather than the underlying sensor signals, achieving 3.29 mm error, a 47 % improvement.
Figure 4: Tasks. We evaluate the performance of the shape-based KineSoft policy across six manipulation tasks. These tasks highlight the advantages of soft robotic hands in contact-rich and delicate object interaction. Red arrows indicate actuation directions for Servo 1; blue arrows indicate actuation directions for Servo 2.
5.3 In-hand Manipulation Task Performance
We evaluated KineSoft on six in-hand manipulation tasks (Figure 4) to showcase both the strengths of soft robots, such as delicate object handling [52, 53, 54] and contact-rich interaction [1, 55], and their conventional challenges, including tasks requiring accurate multi-finger coordination [56].
Bottle Unscrewing:
MOE must rotate and lift a bottle cap using coordinated finger motion.
Lid Flicking:
MOE flicks open a hinged lid on a container.
Berry Picking:
MOE gently picks a berry from a soft branch, evaluating precise grasping and compliant contact with fragile objects inspired by prior works on soft robot fruit harvesting [54].
Paper Grasping:
MOE initiates contact with and lifts a flat sheet of paper from a table.
Fabric Grasping:
MOE conforms around and lifts a loosely draped piece of fabric, highlighting contact-rich manipulation of deformable objects.
Container Unlidding:
MOE grasps and opens the snap-top lid of a plastic container while maintaining a stable grasp. The task highlights multi-finger coordination and localized force application.
Each task was performed in 20 real-world trials using demonstrations collected via kinesthetic teaching, where the operator directly deforms the soft fingers. During execution, KineSoft reproduces these deformations through tendon actuation, bridging the sim-to-real gap with learned shape-aware control. Table 3 reports success rates across all six tasks. KineSoft achieved an 85 % success rate on Bottle Unscrewing, compared to 0 % for a strain-matching baseline based on a state-of-the-art soft in-hand manipulation method [51], implemented with a diffusion policy and trained on the same demonstration data. On Lid Flicking, KineSoft reached 100 % success, while the baseline achieved 90 %. For contact-sensitive tasks such as Berry Picking, Fabric Pickup, and Container Unlidding, KineSoft achieved success rates between 70 % and 80 %, whereas the baseline remained below 35 % on all three. These results highlight the effectiveness of shape-based representations for robust control and reinforce the advantages of soft hands in contact-rich and delicate manipulation settings.
Table 3: Task success rates on 6 different tasks over 20 trials each comparing the baseline strain-matching method to the proposed KineSoft.
Bottle Lid Berry Paper Fabric Container
Method Unscrewing Flicking Picking Grasping Pickup Unlidding
Strain Policy 00/20 18/20 07/20 13/20 01/20 03/20
KineSoft 17/20 20/20 16/20 19/20 13/20 14/20
6 Conclusion and Lessons
This paper presents KineSoft, a framework for learning dexterous manipulation skills with soft robot hands that embraces, rather than fights against, their inherent compliance. The key insight is recognizing that while this compliance enables intuitive kinesthetic teaching, it creates a fundamental gap between demonstration and execution, where the deformations and sensor signals during human demonstration differ substantially from those during autonomous execution through tendon actuation.KineSoft addresses this challenge through a hierarchical approach. A shape estimation model provides consistent geometric representations across demonstration and execution modes while a domain alignment method enables robust transfer of simulation-trained models to real hardware and a shape-conditioned controller reliably tracks the policy’s generated deformation trajectories despite the different underlying actuation mechanisms. Then, a high-level imitation policy learns to generate target vertex deformations from demonstrations, capturing the intended manipulation strategy in geometric grounding. In experiments, results demonstrate that this shape-based hierarchical approach enables more effective skill transfer than methods that attempt to directly match sensor signals or joint configurations. This work suggests that successful imitation learning for dexterous soft robots requires careful consideration of how demonstration and execution modes differ, and appropriate intermediate representations to bridge this gap.
7 Limitations
While KineSoft demonstrates effective shape estimation and trajectory tracking for soft robotic manipulation, several limitations remain. First, the current system lacks explicit force-feedback mechanisms, relying instead on kinematic trajectories and deformation states to encode manipulation skills. This limits precise force control during object interactions. In contrast, human manipulation integrates both proprioception and tactile sensing, suggesting potential benefits from adding complementary tactile input.
Additionally, although kinesthetic teaching provides an robust approach to demonstrating some dexterous tasks, it requires the expert demonstrator to recognize the MOE finger’s workspace limitations to avoid demonstrating shapes that the MOE cannot reach. In future work, we aim to leverage KineSoft’s real-time shape estimation to communicate robot affordances to users via a visual user interface during demonstrations.
Although the sensors used in this work are readily available and low-cost, they introduce fabrication complexity when integrating into the prior MOE design, which was relatively simple to fabricate. We also observed that the sensors may not respond dynamically in fast manipulation tasks due to commonly observed soft body phenomena such as creep and hysteresis, where the sensor signals require some time to reach equilibrium conditions when stretched quickly. For durability, we included results on long-term continuous operation and sensor signal deviation in the Appendix.
KineSoft’s shape-conditioned low-level controller weighs the errors from the vertices of the reference meshes equally. Implicitly, the controller prioritizes the vertices closer to the tip higher because the displacement tends to be greater further away from the base. For many tasks, this is an appropriate weighing as the finger tips play a vital role in performing dexterous in-hand manipulation tasks [57]. However, some tasks may require different regions of the fingers to be prioritized based on contact probability.
Appendix A Task Frames
Figure 5: Tasks
Appendix B Shape Estimation and Tracking Evaluation
Figure 6: Shape estimation and trajectory tracking performance evaluation. We provide each of the shape estimation models and controllers with kinesthetically deformed shape trajectories. A: Shape estimation model comparisons with real-world ground-truth data (red points). B: Shape tracking comparisons with the real-world ground-truth data (red points), references shapes (red), and achieved shapes (blue).
Appendix C Details on Sensor Model
@@
$\frac{L_{i,t}^{R}-L_{i,0}^{R}}{L_{i,0}^{R}}=\frac{1}{1+\frac{\epsilon_{i}}{L_{i,0}^{S}}}\cdot\frac{L_{i,t}^{S}-L_{i,0}^{S}}{L_{i,0}^{S}}.$ (6)
@@
where the observed resistances $R_{i,t},R_{i,0}$ are measured with the DAQ setup. For the $n$ embedded sensors, aligning the simulated and observed distributions involves optimizing the constant correction parameters $\kappa_{0},\kappa_{1},\dots,\kappa_{n-1}$.
@@
Appendix D Baselines
For shape estimation, we compare with analytical and learning-based baselines:
Constant curvature model [49, 41]. Constant curvature model is a common representation for the continuum deformation behavior of soft robot that parametrizes the shape with a single curvature curve [59]. Typically, the independent parameters of the state of the robot are defined by $r_{\mathrm{curve}}$ and $\theta_{\mathrm{curve}}$. Assuming a constant length, $L_{\mathrm{curve}}$ of the robot, we get the constraint:
$L_{\mathrm{curve}}=r_{\mathrm{curve}}\theta_{\mathrm{curve}}.$
In typical applications, additional term $\phi_{\mathrm{curve}}$ is introduced to represent the plane of bending [46]. We implemented this simplified representation for soft robot shape using the proposed strain model as outlined in Section 4.3 and fitting $r_{\mathrm{curve}}$ and $\theta_{\mathrm{curve}}$ to the observed strains in each side of the curve. We transformed the cross-section boundary to the curve during the evaluation and measured the chamfer distance to the reference.